Diffusion morning report 07/07/06

Ted Berg tberg at snf.stanford.edu
Fri Jul 7 06:55:48 PDT 2006


    Good Morning All,
    Looking good this morning . no major problems at this time. Burn box 
being cleaned this morning. Lamp changed in Epi. AMAT working on startup
of ne Epi. Everything else is OK for now.

Diffusion Mike,Ray, Ted 	  	  	  	  	 
  	Date: 	7/7/06 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	Comments 	  	  	 
Tylan 1 	Up 	  	  	  	 
Tylan 2 	Up 	  	  	  	 
Tylan 3 	Up 	  	  	  	 
Tylan 4 	Up 	  	  	  	 
Tylan 5-BBr3 	Up 	  	  	  	 
Tylan 6-POCl3 	Up 	  	  	  	 
Tylan 7-TEOS2 	Up 	  	1.8/30 	  	 
Tylan 8-SiGe 	Up 	  	33.96/30 	  	 
Tylan 9-Poly 	Up 	  	15.33/30 	  	 
Tylan 10-Nitride 	Up 	  	10.13/30 	  	 
Tylan 11-BPSG-LTO 	yellow 	  	34.49/30 	  	 
Tylan 12-FGA 	Up 	  	  	  	 
  	  	  	  	  	 
Thermco 1 	Up 	  	  	  	 
Thermco 2 	Up 	  	  	  	 
Thermco Poly 	Up 	  	  	  	 
Thermco Nitride 	Down 	Testing 	  	  	 
  	  	  	  	  	 
Epi 	Up 	  	  	  	 
  	  	  	  	  	 
Burn Box 	Up 	cleaned this morning 	  	  	 
ZYGO 	up 	  	   	  	 
  	  	  	  	  	 
GAAS 22 	UP 	  	  	  	 
GAAS 23 	up 	  	  	  	 
  	  	  	  	  	 
SiH4 for furnaces 	  	A on line 320 psi 	 changeover working well  	  	 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20060707/6e8ae1d4/attachment.html>


More information about the diffusion mailing list